OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature

OpenStudy (tkhunny):

A few markers will help you see it. -40ºC = -40ºF -- Really Cold 0ºC = 32ºF -- Freezing 40ºC = 104ºF -- Hot enough. 100ºC = 212ºF -- Boiling Those are the ones I always remember for day-to-day usage. You really need only two of them to establish the linear relationship.

OpenStudy (anonymous):

ohh ok and yea I have problems with celcuis a lot

OpenStudy (tkhunny):

You should get \(32 + (9/5)xºC = yºF\) Let's see. \(32 + (9/5)(-40) = 32 - 72 = -40\) ! Yup! That one worked. Boiling. \(32 + (9/5)(100) = 32 + 180 = 212\) ! Yup! That one worked. You do it for "Freeezing".

OpenStudy (anonymous):

lol ok hold on. how did you get 100 then 180

OpenStudy (tkhunny):

100*(9/5) = 9*(100/5) = 9*20 = 180